Given the following information, the Available to Promise will show ________.

Week 1 2 3 4
Forecast Demand 120 120 120 120
Actual Customer Orders 85 100 35 20
Available to Promise
MPS 250 250
  • a. 130, 130, 130, 130
  • b. 165, 65, 30, 230
  • c. 130, 10, -110, 130
  • d. 30, 0, 0, 230
Answer #1

Answer: b. 165.65,30,230

Explanation:

Available to Promise = MPS + Available to Promise for last week - Actual Customer Order

Week 1: APS = 250 + 0 - 85 = 165

Week 2: APS = 0 + 165 - 100 = 65

Week 3: APS = 0 + 65 - 35 = 30

Week 4: APS = 250 + 30 - 20 = 260

Note: In option B, APS for week 4 is incorrect but it is the closest option among all the four options.